# Heterogeneous Multi-treatment Uplift Modeling for Trade-off Optimization in Short-Video Recommendation

### v1

@@ 1. INTRODUCTION
Video recommendation platforms such as YouTube, TikTok, and Kuaishou have become popular in people’s social life for their ability to fulfill users’ diverse and dynamic information needs. While the mainstream recommendation methods focus on designing better ranking models (liu2009learning; ricci2021recommender) (i.e., learning to sort and select the candidate videos that better aligns with the user’s preferences), the industrial solution often engage a series of complementary strategies (or modules) that controls the recommendation statistics to circumvent model biases.
@@
To tackle the task at hand, we propose reformulating the problem using uplift modeling. This personalized causal effect modeling technique is commonly employed in online marketing scenarios, such as distributing coupons and discounts. The primary function of uplift modeling is to measure the incremental impact (i.e., uplift) of treatments or interventions on different populations (efin; flextenet; dragonnet; ganite; cdum), enabling data-driven decision-making. Representative methods include meta-learning models such as T-learner (metalearners) and S-learner (metalearners), as well as tree-based models (wang2015robust; radcliffe2011real; athey2016recursive; nandy2023generalized) and Deep Uplift Modeling (DUM) (he2024rankability; huang2024entire; bnn; euen; liu2024benchmarking; xu2022learning; curth2021nonparametric) . Drawing inspiration from the success of uplift modeling in online marketing, we reformulate the trade-off optimization problem in our video recommendation pipeline as a multi-treatment dynamic uplift modeling task.
@@ 2. RELATED WORK
With the development of deep learning, researchers have found that employing neural networks can introduce more complex and flexible architectures to model the response process to treatment, enabling the learning of more accurate ITE estimates. Therefore, such methods have become a popular framework for uplift modeling (wei2024multi; cevae; bica2020estimating; dragonnet; zhou2023direct). Existing neural network-based Uplift modeling methods can be categorized into two categories: joint modeling of treatment and modeling of specific treatment causal effects. The former estimates the non-treatment score through features and incorporates the treatment feature into the treatment modeling based on this score. For example, EUEN (euen) uses different sub-neural networks to estimate the conversion probability of the control group and the treatment causal effect. EFIN (efin) captures the feature subset related to the control group through explicit feature self-interaction, then models the sensitivity of non-treatment features to specific treatments using a treatment-aware interaction module, and enhances the model’s robustness through an intervention constraint module. M3TN (sun2024m) considers multi-treatment scenarios and uses a multi-task gating network to obtain uplift under different treatments. The latter focuses on modeling the causal effects of specific treatment. DESCN (descn) designs an end-to-end deep cross network to comprehensively capture the information of intervention propensity, response and hidden intervention effects through multi-task learning. CDUM (cdum) expands the treatment representation for guidance and indicator within a multi-task framework and proposes an online real-time feature learning network based on the characteristics of the video recommendation scenario.
@@ 4.1. Offline Hybrid Uplift Modeling
Specifically, take the $k$-th branch as example, we use operations such as multilayer perceptron to generate $M$ expert representations (denoted as $\mathbf{p}_{m}^{k}$, $m\in\{1,2,...,M\}$) from $\mathbf{f}_{i}^{k}$. Following the classical operations of multi-task learning (meng2023parallel; PLE; cigf; copf), we employ a gating mechanism to aggregate the expert information. The final score is then output through a specific tower, using this accumulated information as input. This process can be formulated as:

### v2

@@ 1. INTRODUCTION
Video recommendation platforms such as YouTube, TikTok, and Kuaishou have become popular in people’s social life for their ability to fulfill users’ diverse and dynamic information needs. While the mainstream recommendation methods (cuicalibrating; wu2025aligning) focus on designing better ranking models (liu2009learning; ricci2021recommender; li2024contextual; li2025contrastive; tang2025retrieval) (i.e., learning to sort and select the candidate videos that better aligns with the user’s preferences), the industrial solution often engage a series of complementary strategies (or modules) that controls the recommendation statistics to circumvent model biases.
@@
To tackle the task at hand, we propose reformulating the problem using uplift modeling. This personalized causal effect modeling technique is commonly employed in online marketing scenarios, such as distributing coupons and discounts. The primary function of uplift modeling is to measure the incremental impact (i.e., uplift) of treatments or interventions on different populations (efin; ganite; cdum; li2025klan; sun2024end), enabling data-driven decision-making. Representative methods include meta-learning models such as T-learner (metalearners) and S-learner (metalearners), as well as tree-based models (wang2015robust; radcliffe2011real; athey2016recursive; nandy2023generalized) and Deep Uplift Modeling (DUM) (he2024rankability; huang2024entire; bnn; euen; liu2024benchmarking; xu2022learning; curth2021nonparametric) . Drawing inspiration from the success of uplift modeling in online marketing, we reformulate the trade-off optimization problem in our video recommendation pipeline as a multi-treatment dynamic uplift modeling task.
@@ 2. RELATED WORK
With the development of deep learning, researchers have found that employing neural networks can introduce more complex and flexible architectures to model the response process to treatment, enabling the learning of more accurate ITE estimates. Therefore, such methods have become a popular framework for uplift modeling (wei2024multi; cevae; bica2020estimating; sun2024towards; zhou2023direct). Existing neural network-based Uplift modeling methods can be categorized into two categories: joint modeling of treatment and modeling of specific treatment causal effects. The former estimates the non-treatment score through features and incorporates the treatment feature into the treatment modeling based on this score. For example, EUEN (euen) uses different sub-neural networks to estimate the conversion probability of the control group and the treatment causal effect. EFIN (efin) captures the feature subset related to the control group through explicit feature self-interaction, then models the sensitivity of non-treatment features to specific treatments using a treatment-aware interaction module, and enhances the model’s robustness through an intervention constraint module. M3TN (sun2024m) considers multi-treatment scenarios and uses a multi-task gating network to obtain uplift under different treatments. The latter focuses on modeling the causal effects of specific treatment. DESCN (descn) designs an end-to-end deep cross network to comprehensively capture the information of intervention propensity, response and hidden intervention effects through multi-task learning. CDUM (cdum) expands the treatment representation for guidance and indicator within a multi-task framework and proposes an online real-time feature learning network based on the characteristics of the video recommendation scenario.
@@ 4.1. Offline Hybrid Uplift Modeling
Specifically, take the $k$-th branch as example, we use operations such as multilayer perceptron to generate $M$ expert representations (denoted as $\mathbf{p}_{m}^{k}$, $m\in\{1,2,...,M\}$) from $\mathbf{f}_{i}^{k}$. Following the classical operations of multi-task learning (meng2023parallel; PLE; cigf; copf; tang2024touch; weng2023curriculum), we employ a gating mechanism to aggregate the expert information. The final score is then output through a specific tower, using this accumulated information as input. This process can be formulated as: